Advanced Accounting Ch 1 Q 4

Bobcat, Inc., holds 25 percent of the outstanding shares of Pirate Company and appropriately applies the equity method of accounting.  Excess cost amortization (related to a patent) associated with this investment amounts to $10,000 per year.  For Year 10, Pirate reported earnings of $100,000 and declares and pays cash dividends of $30,000.  During that year, Pirate acquired inventory for $50,000, which it then sold to Bobcat for $80,000.  At the end of Year 10, Bobcat continued to hold merchandise with a transfer price of $32,000. 

 

a.      What Equity in Investee Income should Bobcat report for Year 10?

b.      How will the intra-entity transfer affect Bobcat’s income reporting in Year 11?

c.       If Bobcat had sold the inventory to Pirate, how would the answers to (a) and (b) have changed?
Answer #1

Step 1:  Find gross profit

Sales - Cost of Goods Sold = gross profit

80,000 - 50,000 = 30,000

In % form, 30,000 / 80,000 sales = 37.5%


Step 2: Find deferred Gross profit.

32,000 ending inventory x 37.5% gross profit x 25% portion owned 

= 3,000 deferred gross profit


Step 3: Find investor's portion of net income, amortization.and dividends.

Investor's portion of net Income = 100,000 x 25% = 25,000 net income

Investor's portion of amortization = 10,000

Investor's portion of dividends = 30,000 x 25% = 7,500 (We will ignore dividends in this problem.)


Step 4:


a)  

Find Equity in Earnings 

Equity in Earnings = Portion of Net Income - portion of amortization - deferred gross profit 

25,000 net income - 10,000 amortization - 3,000 deferred gross profit = 12,000 equity in earnings


b)

If the remaining is sold for Year 11, it will increase Year 11's reported net income. If it is not sold, it won't affect it.


c)

Under the equity method, the direction of the sale does not matter.







answered by: anonymous